The diameter of an iron ball used in putting the shot is 12cm.The density of the iron is 7.8g/cm*3.Calculate the mass of the ball in kg to 3significant figures
Answer by ewatrrr(24785)

 
Hi,
.
V = (4/3)pi*r^3 = 288pi
mass =7.8(288pi)g
pi = 3.14